# Bridging the micro-Hz gravitational wave gap via Doppler tracking with the Uranus Orbiter and Probe Mission: Massive black hole binaries, early universe signals and ultra-light dark matter

### v2

@@ Abstract
With the recent announcement by NASA’s Planetary Science and Astrobiology Decadal Survey 2023-2032, a priority flagship mission to the planet Uranus is anticipated. Here, we explore the prospects of using the mission’s radio Doppler tracking equipment to detect gravitational waves (GWs) and other analogous signals related to dark matter (DM) over the duration of its interplanetary cruise. We develop a methodology to stack tracking data and account for time varying detector geometry, thereby constructing the sensitivity of the mission to GWs over the wide frequency range of $3\times 10^{-9}$ Hz to $10^{-4}$ Hz. We find that the mission has the potential to fill the gap between pulsar timing and space-based-interferometry GW observatories. If improvements in reducing Cassini era noise by a factor of 10 are implemented, we forecast the detection of $\mathcal{\mathcal{O}}(\rm{few})$ individual massive black hole binaries using two independent population models. Additionally, we determine the mission’s sensitivity to both astrophysical and primordial stochastic gravitational wave backgrounds, as well as its capacity to test, or even confirm via detection, ultralight DM models. In all these cases, the tracking of the spacecraft over its interplanetary cruise would enable coverage of unexplored regions of parameter space, where signals from new phenomena in our Universe may be lurking.
@@ I Introduction
In this work, we consider how the collection of tracking data over the entirety of the UOP’s interplanetary cruise can increase the sensitivity of the mission to GWs from many sources, allowing us to probe the yet poorly covered micro-Hz regime. Additionally, we outline the mission’s capability to greatly expand the range of ultra-light dark matter (ULDM) models [61, 21] that can be directly detected, including the possibility of a first detection of DM in the Solar System. The key for these results is a method developed here to stack individual short-duration tracking runs over a slowly changing detector geometry, in order to take advantage of the long detector arm constituted by the Earth-spacecraft system. The mission’s GW frequency sensitivity band and DM mass sensitivity are in principle only limited by the total observation time and the data cadence, in analogy to Pulsar Timing Array analyses for nano-Hz GW detection [hereafter PTA, see, e.g., 50, 99, 57, 6, 3].
@@ II.1 Response of Doppler tracking to a passing GW
that are emitted by inspiralling massive black hole binaries (where $f_{\scriptscriptstyle\rm GW}$ is the frequency of the GW, $A$ its amplitude, $\iota$ is the binary’s inclination, and $\phi_{0}$ is an initial phase [see, e.g., 81, 83, 77]); these are the most promising astrophysical source of GWs in the nano-Hz to micro-Hz frequency band, and will be discussed in detail in Section IV.
@@ II.2 Response of Doppler tracking to ultra-light dark matter
If dark matter (DM) is composed of particles of masses $m_{\scriptscriptstyle\rm DM}\lesssim 1$ eV/$c^{2}$ they will be distributed in the Milky Way at inter-distances smaller than their de Broglie wavelength (associated with the ‘size’ of the particle). As a result, DM can be approximated in this range of masses as a classical non-relativistic massive field $\phi_{\scriptscriptstyle\rm DM}$ [61]. From the virialised properties of DM in the Milky Way, this field will be distributed in coherent patches of characteristic length
@@
[…]
@@ II.3 Noise in a Doppler tracking system
Figure 2: Top panel: Observed Cassini PSD for a 40 day run [7] in red, as well as the extrapolation to lower frequencies used in this work (blue, see text). The vertical red bar denotes the representative uncertainty for the low frequency part of the obsered PSD. The gray curve shows the noise PSD based on theoretical work by Ref. [14] and analysis of the Ulysses and Cassini tracking data [30], which was shown to be too conservative. The extrapolation to frequencies lower than $10^{-4}$ Hz [13] is shown as a dashed line. Bottom panel: Response transfer functions to GW for a fixed length Doppler tracking systems with a given incidence angle [111, see also section III]. The strategy to integrate the response over a variable UOP trajectory is detailed in section II.4.
@@ II.4 Strategy for the tracking of low-frequency GWs
Here $f_{\rm{min}}=(10\,{\rm{yr}})^{-1}$ is the inverse of the cruise duration and $f_{\rm{max}}=\tau^{-1}$ is the sampling rate. The coefficients $\mathcal{A}_{i}$ together with the phases $\alpha_{i}$ determine a specific realisation of noise. In the limit of Gaussian noise, the phases are uniformly sampled from $[0,2\pi)$. The amplitudes $\mathcal{A}_{i}$ are instead drawn as random variables from a Gaussian with standard deviation of 1. The frequencies $f_{i}$ are given by all the integer multiples of $f_{\rm{min}}$ up to $f_{\rm max}$, as in a standard Fourier series.
@@
where the tensor sum symbol describes the stacking of several, independent 8 h passes with arbitrary gaps between them. Here $t_{n}$ corresponds to the timestamp at which the $n$-th tracking run is initiated, and $80<t<$ 8 hours follows a clock that is started anew at the beginning of every tracking run. A possible realisation of the virtual tracking data, where the gaps have been removed, is visualised in Fig. 3. We note that from here on, we will assume that the tracking runs are spaced regularly over the mission duration (see Section VII). While this is most likely not realistic, investigating the distribution of tracking runs is only useful once an official mission trajectory is determined. To summarise, in addition to the Allan deviation, each realisation of $y_{2}^{\rm{tot}}$ is thus determined by the parameters characterising both the GW and the spacecraft trajectory. Schematically, they are given by the following:
[…]
@@ III.1 Refresher on sensitivity curves
In this section we estimate how well the GW signal can be extracted from the tracking noise and define a general notion of the mission’s sensitivity curve to GWs that can be coherently modelled, such as those emitted by massive black hole binaries. Typically, the sensitivity curve of a detector to a GW is defined from the detector noise power spectral density $S_{\rm n}(f)$ and the transfer function $R(f)$, which encodes the response of the detector to an incoming GW signal with spectral density $S_{\rm GW}(f)$ [e.g., 71]. The spectral density of the full GW signal plus detector response, per GW cycle is, $S_{\rm GW}(f)R(f)$. Then the optimal SNR of a GW signal is given by [e.g., 96]:
@@ III.2 Noise and response Lomb-Scargle PSDs
The second step is to compute the Lomb Scargle PSDs of the tracking system’s response to monochromatic waves with various frequencies, amplitudes, phases and orientation angles. Figure 5 shows three examples of response PSDs for GW of different frequencies. The time variation in the Doppler tracking system arm length and orientation, as well as the uneven sampling in the tracking runs, produce non-trivial deviations from a delta-function-like response. To define a typical signal, we average the response PSD over the binary inclination angle $\iota$ and the sky location of the source with respect to the ecliptic plane. We highlight here that the response of individual signals will greatly vary from the averaged case, also due to the orbital variation of the UOP.
@@ III.3 Defining the sensitivity to monochromatic GWs
To define a sensitivity curve, we evaluate the integral numerically for monochromatic GW of various frequencies $f$ and subsequently solve for the amplitude $A$, such that SNR$=1$. As a final step, we factor out the number of cycles by reducing the sensitivity with a factor $\sqrt{2fT_{\rm obs}}$. Then we are able to identify the full solution as a function of frequency $A_{\rm sens}^{{\rm{SNR}}=1}(f)$ as the mission’s sensitivity to GWs. For a monochromatic signal with amplitude $A$, the SNR is then simply given by $\sqrt{2fT_{\rm obs}}(A/A_{\rm sens}^{{\rm{SNR}}=1}(f))$, in exact analogy to usual GW wave sensitivity curves. Note that here we must introduce the concept of the mission’s tracking duty cycle $d_{\rm c}$, where:
@@
Figure 6: In the top panel, we compare the GW sensitivity of the prospective Uranus mission to the current sensitivity of PTAs [37] and the sensitivity of the recently adopted LISA mission. The three black/grey lines correspond to the scenarios detailed in Section III. In the bottom panel, we show a waterfall plot of the mission’s hypothetical detection horizon for chirping equal-mass (here masses are in the observer’s frame) massive black hole binaries, assuming a maximum of 10 years for the signal duration and the priority scenario, as described in Section III.5. We compare it to the same plot for PTAs and for LISA (assuming 4 years of maximal signal duration). Note that monochromatic sources, expected for PTAs and from our population models in §IV, will have lower SNR for the same mass and redshift in the bottom panel.
@@ III.5 Baseline, Priority and Breakthrough scenarios
The monochromatic sensitivity curves constructed above (Fig. 6) span the interesting frequency range of few $\sim 10^{-9}$ Hz to $\sim 10^{-4}$ Hz. Below $\sim 10^{-6}$ Hz, the slope is approximately $f^{-1}$, where the sensitivity is determined by the extrapolation of the Cassini noise. At higher frequencies, the slope mellows to approximately $\sim f^{-1/2}$ before a turning point at $f\sim$1/(8 hours) where sensitivity is additionally affected by the discretisation of the observation period. We truncate it at few ×$\times 10^{-4}$ Hz to assure that numerical noise due to the Nyquist frequency is completely absent. The features of the sensitivity are a result of the interaction between the GW response and the time changing detector geometry, the un-even sampling of the tracking data as well as inheriting the noisiness of the observed Cassini PSD. Crucially, strain sensitivities of the order few $\sim 10^{-14}$ in the micro-Hz band can be achieved without requiring significant improvements in the Allan deviation, provided that the mission’s duty cycle is of order $\sqrt{d_{\rm c}}\sim 1$, i.e., that tracking data is accumulated over a substantial amount of the entire mission duration. We note that this amount of tracking is not typical for deep interplanetary missions. Here, however, we argue that it should be prioritised in parallel with Allan deviation improvements in order to achieve the scientific objectives detailed in this manuscript. We assume this level of commitment to the tracking and fix a duty cycle of $d_{\rm c}=1/3$, which arises from a total of approximately 3650, 8-hour tracking runs. We note here that switching to multiple tracking stations to achieve $d_{\rm c}>1/3$ should not, in principle, introduce additional sources of noise as long as the phase of the GW is not lost due to a lack of clock synchronization between the stations. Considering the low frequency of the tracked GWs, this is not an issue here.
@@ IV Detection forecast for Black Hole binaries
Supermassive black hole binaries (SMBHBs) are among the loudest sources in the low-frequency GW sky. They are expected as a consequence of galaxy mergers, although how frequently they form and how they evolve in their environments is still not fully understood [see, e.g., the seminal papers 12, 115, 108, 59, 69]. Recent constraints on a GW background signal by pulsar timing arrays (PTA) provide evidence of an SMBHB population in the mass range $\sim 10^{8}-10^{10}M_{\odot}$ [3, 6, 94, 119] while the direct detections of lighter SMBHB mergers below a few $10^{7}M_{\odot}$ is expected in the 2030s with the advent of space-based interferometers such as LISA [29] and TianQin [75]. As we will show, the sensitivity derived in Section III facilitates the detection of SMBHBs that are evolving beyond the nano-Hz band and into the milli-Hz regime. As an illustration, we show the hypothetical sensitivity of the mission to chirping, equal-mass BH binaries as a waterfall plot in the bottom panel of Fig. 6. The mission’s peak sensitivity lies in the total mass range of $10^{7}$ to $10^{10}M_{\odot}$. Such heavy, chirping binaries would in principle be detectable with high confidence up to redshifts of $z\sim 3$ for an Allan deviation of $\sigma_{\rm A}=10^{-14}$ and essentially over the whole cosmological volume for $\sigma_{\rm A}=1\times 10^{-15}$.
@@ IV.1 Model 1: Millennium simulation
×$\times\frac{{\rm d}n_{\bullet}}{{\rm d}M_{\bullet}}\frac{{\rm d}^{2}\Gamma}{{\rm d}\xi{\rm d}z}(\xi,z_{\rm{del}}),$ (30)
@@
×$\times{\rm d}f_{\scriptscriptstyle\rm{GW}}\,{\rm d}\xi{\rm d}M_{\bullet}{\rm d}z,$
@@
Pop. SNR Nr. Det Millennium 1 0.2×$\times P_{\rm occ}$ Millennium 3 $0.02\times P_{\rm occ}$ Millennium 8 0           Pop. SNR Nr. Det Millennium 1 1.2×$\times P_{\rm occ}$ Millennium 3 0.2×$\times P_{\rm occ}$ Millennium 8 0.03×$\times P_{\rm occ}$           Pop. SNR Nr. Det Millennium 1 24.2×$\times P_{\rm occ}$ Millennium 3 5.9×$\times P_{\rm occ}$ Millennium 8 1.6×$\times P_{\rm occ}$
@@
Pop. SNR Nr. Det Millennium 1 0.08×$\times P_{\rm occ}$ Millennium 3 $0.009\times P_{\rm occ}$ Millennium 8 0 holodeck 1 $0.02^{+0.01}_{-0.002}$ holodeck 3 00 holodeck 8 00        Pop. SNR Nr. Det Millennium 1 0.5×$\times P_{\rm occ}$ Millennium 3 0.08×$\times P_{\rm occ}$ Millennium 8 0.01×$\times P_{\rm occ}$ holodeck 1 $0.03^{+0.1}_{0.02}$ holodeck 3 00 holodeck 8 00           Pop. SNR Nr. Det Millennium 1 14.6×$\times P_{\rm occ}$ Millennium 3 3.1×$\times P_{\rm occ}$ Millennium 8 0.7×$\times P_{\rm occ}$ holodeck 1 $0.8^{+2.9}_{-0.7}$ holodeck 3 $0.11^{+0.4}_{-0.1}$ holodeck 8 $0.015^{+0.04}_{-0.011}$

### v3

@@ Abstract
With the recent announcement by NASA’s Planetary Science and Astrobiology Decadal Survey 2023-2032, a priority flagship mission to the planet Uranus is anticipated. Here, we explore the prospects of using the mission’s radio Doppler tracking equipment to detect gravitational waves (GWs) and other analogous signals related to dark matter (DM) over the duration of its interplanetary cruise. We develop a methodology to stack tracking data and account for time varying detector geometry, thereby constructing the sensitivity of the mission to GWs over the wide frequency range of $3\times 10^{-9}$ Hz to $10^{-4}$ Hz. We find that the mission has the potential to fill the gap between pulsar timing and space-based-interferometry GW observatories. If improvements in reducing Cassini era noise by a factor of $\mathcal{O}(10)$ are implemented, we forecast the detection of $\mathcal{\mathcal{O}}(\rm{few})$ individual massive black hole binaries using two independent population models. Additionally, we determine the mission’s sensitivity to both astrophysical and primordial stochastic gravitational wave backgrounds, as well as its capacity to test, or even confirm via detection, ultralight DM models. In all these cases, the tracking of the spacecraft over its interplanetary cruise would enable coverage of unexplored regions of parameter space, where signals from new phenomena in our Universe may be lurking.
@@ I Introduction
In this work, we consider how the collection of tracking data over the entirety of the UOP’s interplanetary cruise can increase the sensitivity of the mission to GWs from many sources, allowing us to probe the yet poorly covered micro-Hz regime. Additionally, we outline the mission’s capability to greatly expand the range of ultra-light dark matter (ULDM) models [62, 21] that can be directly detected, including the possibility of a first detection of DM in the Solar System. The key for these results is a method developed here to stack individual short-duration tracking runs over a slowly changing detector geometry, in order to take advantage of the long detector arm constituted by the Earth-spacecraft system. The mission’s GW frequency sensitivity band and DM mass sensitivity are in principle only limited by the total observation time and the data cadence, in analogy to Pulsar Timing Array analyses for nano-Hz GW detection [hereafter PTA, see, e.g., 51, 100, 58, 6, 3].
@@ II.1 Response of Doppler tracking to a passing GW
that are emitted by inspiralling massive black hole binaries (where $f_{\scriptscriptstyle\rm GW}$ is the frequency of the GW, $A$ its amplitude, $\iota$ is the binary’s inclination, and $\phi_{0}$ is an initial phase [see, e.g., 82, 84, 78]); these are the most promising astrophysical source of GWs in the nano-Hz to micro-Hz frequency band, and will be discussed in detail in Section IV.
@@ II.2 Response of Doppler tracking to ultra-light dark matter
[…]
@@
$\begin{split}\psi(t)\approx\psi_{g}(t)&\equiv\,6\cdot 10^{-16}\left(\frac{\rho_{\scriptscriptstyle\rm DM}}{0.3\,\mathrm{GeV}/\mathrm{cm}^{3}}\right)\\ &\left(\frac{10^{-23}\,\mathrm{eV}/c^{2}}{m_{\scriptscriptstyle\rm DM}}\right)^{2}\cos(4\pi f_{\scriptscriptstyle\rm DM}t+2\alpha_{0}),\end{split}$ (9)
@@ II.3 Noise in a Doppler tracking system
Figure 2: Top panel: Observed Cassini PSD for a 40 day run [7] in red, as well as the extrapolation to lower frequencies used in this work (blue, see text). The vertical red bar denotes the representative uncertainty for the low frequency part of the obsered PSD. The gray curve shows the noise PSD based on theoretical work by Ref. [14] and analysis of the Ulysses and Cassini tracking data [30], which was shown to be too conservative. The extrapolation to frequencies lower than $10^{-4}$ Hz [13] is shown as a dashed line. Bottom panel: Response transfer functions to GW for a fixed length Doppler tracking systems with a given incidence angle [112, see also section III]. The strategy to integrate the response over a variable UOP trajectory is detailed in section II.4.
@@ II.4 Strategy for the tracking of low-frequency GWs
Here $f_{\rm{min}}=(10\,{\rm{yr}})^{-1}$ is the inverse of the cruise duration and $f_{\rm{max}}=\tau^{-1}$ is the sampling rate. The coefficients $\mathcal{A}_{i}$ together with the phases $\alpha_{i}$ determine a specific realisation of noise. In the limit of Gaussian noise, the phases are uniformly sampled from $0,2\pi)$. The amplitudes $\mathcal{A}_{i}$ are instead drawn as random variables from a Gaussian with standard deviation of 1. The frequencies $f_{i}$ are given by all the integer multiples of $f_{\rm{min}}$ up to $f_{\rm max}$, as in a standard Fourier series.
@@
where the tensor sum symbol describes the stacking of several, independent 8 h passes with arbitrary gaps between them. Here $t_{n}$ corresponds to the timestamp at which the $n$-th tracking run is initiated, and $0<t<8$ hours follows a clock that is started anew at the beginning of every tracking run. A possible realisation of the virtual tracking data, where the gaps have been removed, is visualised in Fig. 3. We note that from here on, we will assume that the tracking runs are spaced regularly over the mission duration (see Section VII). While this is most likely not realistic, investigating the distribution of tracking runs is only useful once an official mission trajectory is determined. To summarise, in addition to the Allan deviation, each realisation of $y_{2}^{\rm{tot}}$ is thus determined by the parameters characterising both the GW and the spacecraft trajectory. Schematically, they are given by the following:
GW $\text${GW $\to$ $\begin{cases}\hat{\bm{k}};&\text{Propagation vector}\\ f_{\scriptscriptstyle\rm GW};&\text{Frequency}\\ A;&\text{Amplitude}\\ \phi_{0};&\text{Phase}\end{cases}$
$\text{Trajectory}\to$ $\begin{cases}\hat{\bm{n}}(t_{n};t);&\text{Link unit vector}\\ T_{2}(t_{n};t);&\text{Two-way travel time}\end{cases}$
@@ III.1 Refresher on sensitivity curves
In this section we estimate how well the GW signal can be extracted from the tracking noise and define a general notion of the mission’s sensitivity curve to GWs that can be coherently modelled, such as those emitted by massive black hole binaries. Typically, the sensitivity curve of a detector to a GW is defined from the detector noise power spectral density $S_{\rm n}(f)$ and the transfer function $R(f)$, which encodes the response of the detector to an incoming GW signal with spectral density $S_{\rm GW}(f)$ [e.g., 72]. The spectral density of the full GW signal plus detector response, per GW cycle is, $S_{\rm GW}(f)R(f)$. Then the optimal SNR of a GW signal is given by [e.g., 97]:
@@ III.2 Noise and response Lomb-Scargle PSDs
The second step is to compute the Lomb Scargle PSDs of the tracking system’s response to monochromatic waves with various frequencies, amplitudes, phases and orientation angles. Figure 5 shows three examples of response PSDs for GW of different frequencies. The time variation in the Doppler tracking system arm length and orientation, as well as the uneven sampling in the tracking runs, produce non-trivial deviations from a delta-function-like response. To define a typical signal, we average the response PSD over the binary inclination angle $\iota$ and the sky location of the source with respect to the ecliptic plane. We highlight here that the response of individual signals will greatly vary from the averaged case. Interestingly, the mission lacks ”blindspots” in for any longitudinal angle due to the presence of curvature in the UOP trajetcory (See Appendix B for a plot of the expected SNR as a function of source angles). Otherwise, the averaging procedure is analogous to standard practive in GW detectors [see e.g. 97]
@@ III.3 Defining the sensitivity to monochromatic GWs
To define a sensitivity curve, we evaluate the integral numerically for monochromatic GW of various frequencies $f$ and subsequently solve for the amplitude $A$, such that SNR=1=1. As a final step, we factor out the number of cycles by reducing the sensitivity with a factor $\sqrt{2fT_{\rm obs}}$. Then we are able to identify the full solution as a function of frequency $A_{\rm sens}^{{\rm{SNR}}=1}(f)$ as the mission’s sensitivity to GWs. For a monochromatic signal with amplitude $A$, the SNR is then simply given by $\sqrt{2fT_{\rm obs}}(A/A_{\rm sens}^{{\rm{SNR}}=1}(f))$, in exact analogy to usual GW wave sensitivity curves. Note that here we must introduce the concept of the mission’s tracking duty cycle $d_{\rm c}$, where:
@@
Figure 6: In the top panel, we compare the GW sensitivity of the prospective Uranus mission to the current sensitivity of PTAs [37] and the sensitivity of the recently adopted LISA mission. The three black/grey lines correspond to the scenarios detailed in Section III. In the bottom panel, we show a waterfall plot of the mission’s hypothetical detection horizon for chirping equal-mass (here masses are in the observer’s frame) massive black hole binaries, assuming a maximum of 10 years for the signal duration and the priority scenario, as described in Section III.5. We compare it to the same plot for PTAs and for LISA (assuming 4 years of maximal signal duration). The SNR iso-lines start at SNR=1 and are denote factor 10 increases in SNR. Note that monochromatic sources, expected for PTAs and from our population models in §IV, will have lower SNR for the same mass and redshift in the bottom panel.
@@ III.5 Baseline, Priority and Breakthrough scenarios
The monochromatic sensitivity curves constructed above (Fig. 6) span the interesting frequency range of few $\sim 10^{-9}$ Hz to $\sim 10^{-4}$ Hz. Below $\sim 10^{-6}$ Hz, the slope is approximately $f^{-1}$, where the sensitivity is determined by the extrapolation of the Cassini noise. At higher frequencies, the slope mellows to approximately $\sim f^{-1/2}$ before a turning point at $f\sim$1/(8 hours) where sensitivity is additionally affected by the discretisation of the observation period. We truncate it at few $\times 10^{-4}$ Hz to assure that numerical noise due to the Nyquist frequency is completely absent. The features of the sensitivity are a result of the interaction between the GW response and the time changing detector geometry, the un-even sampling of the tracking data as well as inheriting the noisiness of the observed Cassini PSD. Crucially, strain sensitivities of the order few $\sim 10^{-14}$ in the micro-Hz band can be achieved without requiring significant improvements in the Allan deviation, provided that the mission’s duty cycle is of order $\sqrt{d_{\rm c}}\sim 1$, i.e., that tracking data is accumulated over a substantial amount of the entire mission duration. We note that this amount of tracking is not typical for deep interplanetary missions. Here, however, we argue that it should be prioritised in parallel with Allan deviation improvements in order to achieve the scientific objectives detailed in this manuscript. We assume this level of commitment to the tracking and fix a duty cycle of $d_{\rm c}=1/3$, which arises from a total of approximately 3650, 8-hour tracking runs. We note here that switching to multiple tracking stations to achieve $d_{\rm c}>1/3$ should not, in principle, introduce additional sources of noise as long as the phase of the GW is not lost due to a lack of clock synchronization between the stations. Considering the low frequency of the tracked GWs, this is not an issue here.
@@ IV Detection forecast for Black Hole binaries
Supermassive black hole binaries (SMBHBs) are among the loudest sources in the low-frequency GW sky. They are expected as a consequence of galaxy mergers, although how frequently they form and how they evolve in their environments is still not fully understood [see, e.g., the seminal papers 12, 116, 109, 60, 70]. Recent constraints on a GW background signal by pulsar timing arrays (PTA) provide evidence of an SMBHB population in the mass range $\sim 10^{8}-10^{10}M_{\odot}$ [3, 6, 95, 120] while the direct detections of lighter SMBHB mergers below a few $10^{7}M_{\odot}$ is expected in the 2030s with the advent of space-based interferometers such as LISA [29] and TianQin [76]. As we will show, the sensitivity derived in Section III facilitates the detection of SMBHBs that are evolving beyond the nano-Hz band and into the milli-Hz regime. As an illustration, we show the hypothetical sensitivity of the mission to chirping, equal-mass BH binaries as a waterfall plot in the bottom panel of Fig. 6. The mission’s peak sensitivity lies in the total mass range of $10^{7}$ to $10^{10}M_{\odot}$. Such heavy, chirping binaries would in principle be detectable with high confidence up to redshifts of $z\sim 3$ for an Allan deviation of $\sigma_{\rm A}=10^{-14}$ and essentially over the whole cosmological volume for $\sigma_{\rm A}=1\times 10^{-15}$.
@@ IV.1 Model 1: Millennium simulation
$\times\frac{{\rm d}n_{\bullet}}{{\rm d}M_{\bullet}}\frac{{\rm d}^{2}\Gamma}{{\rm d}\xi{\rm d}z}(\xi,z_{\rm{del}}),$ (30)
@@
$\times{\rm d}f_{\scriptscriptstyle\rm{GW}}\,{\rm d}\xi{\rm d}M_{\bullet}{\rm d}z,$
@@
[…]
@@
[…]